OpenStudy (anonymous):

it is a blistering -hot day at the beach .You look at your portable thermometer , which says it is 35 c . what is the temperature in fahrenheit and kevin ? show the equations youy use to make these conversions . it is a blistering -hot day at the beach .You look at your portable thermometer , which says it is 35 c . what is the temperature in fahrenheit and kevin ? show the equations youy use to make these conversions . @Physics

OpenStudy (agreene):

so, the easiest to get is Kelvin, because: \[K=C+273.15\] Where K is degrees Kelvin and C is degrees Centigrade K=35+273.15 K=308.15 C to F is a bit harder \[F=(C-1.8)*32\] So, F=(35-1.8)*32 F=1062.4